Circular No. 6014

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET McNAUGHT-HARTLEY (1994n) R. H. McNaught reports his discovery of a comet on a 110-min ex-

posure on July 6 by M. Hartley with the 1.2-m U.K. Schmidt telescope at Siding Spring, and he provides the following measurements, also from a plate obtained the night before in very bad seeing. The discovery plate shows a very strong condensation, a 15′′ coma and a tail in p.a. 270^{o}.

1994 UT  α_{2000}  δ_{2000}  m_{1}
July 5.67164  21^{h}36^{m}25.^{s}66  −41^{o}43′55′′.6
6.60581  21 36 19.76  −41 50 17.0  16
6.68220  21 36 19.03  −41 50 47.9

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Precise positions measured by D. D. Balam from CCD images by J. B.

Tatum and himself with the 0.5-m reflector at the Climenhaga Observatory:

1994 UT  α_{2000}  δ_{2000}
July 7.37549  3^{h}53^{m}06.^{s}04  +70^{o}10′55′′.5
7.38414  3 53 04.73  +70 10 54.2
7.39098  3 53 03.74  +70 10 54.1

SUPERNOVA 1993J IN NGC 3031 H.-U. Zimmermann, B. Aschenbach, G. Hasinger, W. Pietsch, P. Pre-

dehl and J. Trümper, Max-Planck-Institut für Extraterrestrische Physik; W. Lewin and E. Magnier, Massachusetts Institute of Technology; J. van Paradijs, University of Amsterdam; G. Fabbiano, Harvard-Smithsonian Center for Astrophysics; L. Lubin, Princeton University; and R. Petre, NASA Goddard Space Flight Center, report: “One year after the original detection, supernova 1993J was reobserved by the x-ray observatory ROSAT. A 2000-s exposure taken on 1994 Apr. 4 showed the source at a rate of 0.018 ± 0.003 cts/s in the energy range 0.1–2.4 keV of the PSPC detector. With the PSPC measurements from early May 1993 (when the soft x-ray rate had settled at 0.052 ± 0.002 cts/s after the rapid decrease during the first few weeks) and Nov. 1993 (0.035 ± 0.001 cts/s), the new observation suggests that the x-ray count rate has decreased linearly over the past year with a gradient of ∼ 0.038 cts s^{−1} yr^{−1}. As reported previously by us (IAUC 5899), the temperature decreased significantly during the first half year, while the absorbing column density increased. The new data continue to show this trend.”

1994 July 7 Brian G. Marsden